OpenStudy (anonymous):

Use the compound interest formula P = A(1 + i)n, where A is the original value of an investment, i is the interest rate per compounding period, n is the total number of compounding periods, and P is the value of the investment after n periods. A financial advisor recommends that a client deposit $2900 into a fund that earns 7.5% annual interest compounded monthly. What will be the value of the investment after 6 years? Round to the nearest cent.

hero (hero):

P = 2900(1+.075)^6

hero (hero):

P = 4475.57

OpenStudy (anonymous):

that answer is right

OpenStudy (anonymous):

not*

hero (hero):

Okay sorry

OpenStudy (anonymous):

it oaky just want to understand it

hero (hero):

P = 2900(1+.075)^72 12*6 = n 72 equal the total number of months compounded, so the correct answer is P = 529429

hero (hero):

I'm not going to argue with my calculator

hero (hero):

It compounds exponentially so that appears to be correct

OpenStudy (anonymous):

ya it is

OpenStudy (anonymous):

Use the exponential decay equation given below, where A is the amount of a radioactive material present after time t, k is the half-life of the radioactive material, and A0 is the original amount of radioactive substance. Iodine-131 is an isotope that is used to study the functioning of the thyroid gland. This isotope has a half-life of approximately 8 days. A patient is given an injection that contains 8 micrograms of iodine-131. (Round your answers to one decimal place.) (a) What will be the amount of iodine in the patient after 15 days? A = micrograms (b) How long (in days) will it take for the amount of iodine in the patient to reach 5 micrograms? t = days

hero (hero):

I don't see the radio active decay formula

OpenStudy (anonymous):

srry